OpenStudy (anonymous):

Given the linear equation y= -3/7x-1, find the y- coordinates of the points (-7,),(0,), and (7,). Please show all work.

OpenStudy (anonymous):

y=(-3/7)(-7)-1=3-1=2 y=(-3/7)(0)-1=-1 y=(-3/7)(7)-1=-3-1=-4